OpenStudy (anonymous):

If 6 newborn babies are randomly selected, how many different gender sequences are possible?

OpenStudy (anonymous):

Is this for math or science? Since you have 2 possible birth genders and 6 babies, I'd say...12? 6/12 = 1/2...I'm really not sure about this :(

OpenStudy (anonymous):

Its for math combinations.

OpenStudy (anonymous):

Answer choices?

OpenStudy (anonymous):

a. 720 c. 12 b. 36 d. 64

OpenStudy (anonymous):

The answer turned out to be 64 when i went back and redid my math thanks

OpenStudy (anonymous):

I'd probably go with C...However, I suggest getting a second opinion, as I'm not sure what level of math you're in and if that is the process you would use to get through that problem. What steps have you used on similar ones you've gotten right?

OpenStudy (anonymous):

How did you get 64? (Not that I doubt it was right, I'm just curious how you did this).

OpenStudy (mathmate):

There are two choices of gender for each child, and there are 6 children, so the number of arrangements is \[2^6=64\]

OpenStudy (mathmate):

@shopgirl1701

OpenStudy (anonymous):

Interesting. Thanks @mathmate :)

OpenStudy (mathmate):

You're welcome! :)